# Dispersion on Time-Varying Graphs

### v1

@@ 2 Impossibility Results
(Impossibility without global communication) It is impossible to solve Dispersion on TVG with $k=n$ agents for any $n\geq 6$ with the agents having $1$-hop visibility but no global communication even if agents are initially co-located and have infinite memory.
@@
We prove the above theorem by contradiction. Let us consider a ring $\mathcal{R}$ of size $n$ and $n$ many agents be co-located at some node of the ring $\mathcal{R}$. Suppose an algorithm $\mathcal{A}$ solves Dispersion on TVG. If agents are not in the dispersed configuration at the beginning of round $r$, it implies that there is at least one multinode at the beginning of round $r$. Our idea is to show that there must exist a multinode in the beginning round $r+1$. Observe that, if there is some multinode of size at least 4 at the beginning of round $r$ then there must exist a multinode at the beginning of the next round. Based on the size of the multinode, we categorise this into two cases.
  • Case 1: At the beginning round $r$, there is at least a multinode of size 3.
  • Case 2: At the beginning round $r$, there is at least one multinode of size 2 and there is no multinode of size more than 2.
Let $v$ be multinode in the ring $\mathcal{R}$ at the beginning round $r$, and $v_{1}$ and $v_{2}$ be two neighbours of node $v$ in the ring $\mathcal{R}$. Since $n\geq 6$, except node $v$, let $w_{1}$ and $w_{2}$ be neighbours of $v_{1}$ and $v_{2}$ respectively on the footprint ring. Agents at node $w_{1}$ can not see the information of node $v_{2}$, $w_{2}$ and $v$ using 1-hop visibility. Similarly, agents at node $w_{2}$ can not see the information of node $v_{1}$, $w_{1}$ and $v$. Therefore, according to $\mathcal{A}$, the decision of the move cycle in each round of agents in node $v_{1}$ is independent of the decision of the move cycle of agents at node $v_{2}.$ At the start of round $r$, if the adversary removes an edge according to the following strategy then at least a multinode remains at the start of the next round.
  • Strategy for case 1: Let $v$ be a multinode of size 3. At the starting of round $r$, the adversary removes edge $(v,\,v_{1})$. In this case, no matter what steps agents follow, there will be a multinode of the size of at least two in the next round.
  • Strategy for case 2: Let $v$ be a multinode of size two. We divide this into three cases, depending on the situation at the beginning of round $r$.
    • Both $v_{1}$ and $v_{2}$ are holes
      If in round $r$, no agent moves to the node $v_{1}$ and $v_{2}$ from their neighbours except node $v$, then there is always a hole present at the start of round $r+1$.
      Else, if in round $r$, agent(s) move to the node $v_{1}$ and $v_{2}$ from their neighbours $w_{1}$ and $w_{2}$ respectively, then there is always a multinode present among $v_{1},v,v_{2}$ at the start of the next round.
      Else, in round $r$, agent(s) move to $v_{1}$ (or $v_{2}$) from its neighbour $w_{1}$ (or $w_{2}$). W.l.o.g., agent(s) decides to move from node $w_{1}$ to node $v_{1}$ in round $r$. In this case, the adversary removes the edge $(v,\,v_{2})$ at the start of round $r$. At the round $r$, agents from node $v$ can not move to the node $v_{2}$ due to the missing edge $(v,\,v_{2})$. They either can stay at node $v$ or move to the node $v_{1}$. In both cases, there is a multinode at the start of the next round. Specifically, if agent(s) at node $v$ decides to move, it creates a multinode at node $v_{1}$; otherwise, node $v$ remains as a multinode at the start of round $r+1$.
    • Either $v_{1}$ or $v_{2}$ is a hole W.l.o.g. let $v_{1}$ be a hole.
      If in round $r$, one agent moves from both neighbours of node $v_{1}$, the adversary does not remove any edge at the start of round $r$. In this case, it creates a multinode at node $v_{1}$ at the start of round $r+1$.
      Else, if in round $r$, agent(s) from one of the nodes $v$, $w_{1}$, say from $w$, decides to move to the node $v_{1}$, the adversary removes edge $(v_{1},\,w)$ at the start of round $r$. In this case, node $v_{1}$ remains a hole at the start of the next round.
      Otherwise, in round $r$, no agent moves to the node $v_{1}$ in the round $r$, then node $v_{1}$ remains hole at the start of round $r+1$.
    • Neither $v_{1}$ nor $v_{2}$ is a hole
      If in round $r$, agents at node $v_{1}$ and $v_{2}$ do not move in round $r$, the adversary does not remove any edge. In this case, one among $v_{1},v,v_{2}$ is a multinode at the start of the next round.
      Otherwise, if in round $r$, agent(s) from node $v_{1}$ and $v_{2}$ move to node $v$, then the adversary does not remove any edge. In this case, there will be a multinode at the node $v$.
      Else, if in round $r$, agent(s) of either $v_{1}$ or $v_{2}$, w.l.o.g. say of $v_{1}$, does not move in round $r$, then the adversary removes the edge $(v,\,v_{2})$. In this case, there will be a multinode at $v$ or $v_{1}$ at the start of the next round.
      Otherwise, if in round $r$, agent(s) from node $v_{1}$ and $v_{2}$ move to node $w_{1}$ and $w_{2}$, respectively, the adversary does not remove any edge. In this case, no matter how agents move from node $v$, it creates a hole or a multinode at the start of round $r+1$.
      Else, in round $r$, agent(s) of either $v_{1}$ or $v_{2}$ move to $v$, say agent(s) of $v_{1}$ move to node $v$. In this case, if the agent from node $v_{2}$ also moves to node $v$, then the adversary does not remove any edge. At the beginning of the next round $r+1$, there is a multinode at node $v$. If the agent at $v_{2}$ decides to stay at node $v_{2}$ in round $r$, then also the adversary does not remove any edge. In this case, there is a multinode at node $v$ or $v_{2}$ at the beginning of the next round. If agent from $v_{2}$ moves toward to $w_{2}$, then the adversary deletes edge $(v_{2},\,w_{2})$. In this case, no matter how the agents move from node $v_{2}$, there is a multinode at node $v_{2}$ or $v$ at the beginning of round $r+1$.
@@
(Impossibility without 1-hop visibility) It is impossible to solve Dispersion on TVG with $k=n$ agents considering global communication but without $1$-hop visibility even if the agents are initially co-located and have infinite memory.
@@
If the agents have zero visibility, then the choice of movement of each agent depends only on the information present at the current node and the information attained via global communication. Since the algorithm is deterministic, the adversary can remove an edge from $G$ based on the movement of the agents. We prove the above theorem by showing the impossibility of achieving dispersion on rings. Let the agents follow algorithm $\mathcal{A}$ to achieve dispersion. In a given round, if the movement of agents does not lead to dispersion, the adversary does not need to do anything. If the movement of agents leads to dispersion, the adversary has to intervene. We have the following possible scenarios:
  • If there is a hole, say $v$, in the ring and the agents from both the neighbouring nodes of nodes $v$ are going to enter into the node: In this case, the adversary does not need to do anything as a new multinode is created in the ring.
  • If there is a hole in the ring and the agent from exactly one neighbouring node is going to enter into the hole: In this case, the adversary removes the edge connecting the hole with this neighbouring node (recall that in our model agents at a node $v$ can not sense whether an edge associated with node $v$ is missing or not). Therefore, the movement of the agent becomes unsuccessful in this round. Thus, the hole remains, and dispersion is not achieved.
The adversary can prevent the agents from achieving dispersion in any given round $r$ using the above-mentioned strategies. ∎
Remark 1. From [7], the time lower bound is $\Omega(k)$ to solve dispersion on static graphs for $k\leq n$. From [2], the memory lower bound is $\Omega(\log k)$ per agent to solve dispersion on static graphs for $k\leq n$. We are working on a dynamic graph model which is weaker than a static graph model. Therefore, the lower bound for time and memory per agent to solve Dispersion on TVG for $k\leq n$ is $\Omega(k)$ and $\Omega(\log k)$, respectively. Due to Theorem 2.1 and 2.2 to solve Dispersion on TVG $\mathcal{G}$ for $k\leq n$, agents need 1-hop visibility and global communication. In [10], when agents are equipped with 1-hop visibility and global communication, authors provide an algorithm to solve dispersion in $O(k)$ time and $O(\log k)$ memory per agent in their dynamic graph model. Therefore, we can use the same algorithm and solve Dispersion on TVG for $k\leq n$ agents in $\Theta(k)$ time and $\Theta(\log k)$ memory per agent as TVG exhibits a stronger graph model.
@@ 3 The Dispersion on General Graphs
In this section, we introduce an algorithm that solves the Dispersion on TVG when there are $n+1$ agents co-located at some node in a general graph. We use the idea of depth-first search (DFS) traversal by mobile agents that also solves the dispersion on static graphs [2]. For the sake of completeness, let us recall the idea of DFS traversal. At each node $v$, an agent chooses an unexplored port $p$ and traverses through it: (i) if it reaches an unvisited node $u$ (agents can understand this by the fact that there is no agent settled at $u$), then one agent settles at $u$ and stores port $p^{\prime}$ (through which it entered node $u$) as its parent port. Now, the unsettled agent tries to move through port $(p^{\prime}+1)$ mod $\delta_{u}$ in the explore state, (ii) else, if the agent reaches an already visited node (it can understand this because there is a settled agent present at that node), the agent returns to node $v$ in the backtrack state. If an unsettled agent reaches node $v$ in the backtrack state, then it computes $q=(p+1)$ mod $\delta_{v}$. If port $q$ is unexplored, it moves through port $q$ in the explore state. Else, if port $q$ equals the parent port of $v$, then it moves through port $q$ in the backtrack state. This algorithm takes $4m$ rounds to run on any graph, where $m$ is the number of edges in that graph. The information required for an agent at each node to run this algorithm can be stored by the settled agent that is already present at that node.
@@
High-level idea: No group of agents moves indefinitely through which its movement was unsuccessful. Initially, there is only one group as the configuration is rooted. Whenever the group has an unsuccessful move, the agents divide into two groups, namely $G_{1}$ and $G_{2}$. These two groups run their dispersion algorithm via DFS separately. A missing edge can make the movement unsuccessful for both groups. On unsuccessful movement, $G_{1}$ never deviates from its DFS traversal path, whereas $G_{2}$ deviates after some finite waiting period. This keeps the algorithm running and either $G_{2}$ gets dispersed or $G_{1}$ can move one hop according to its DFS traversal. If $G_{2}$ gets dispersed, $G_{1}$ can understand it within some finite rounds and gets divided into two groups further. Else, if $G_{1}$ gets dispersed by moving one hop again and again, $G_{2}$ can understand it within some finite rounds and gets divided into two groups further. We ensure that none of $G_{1}$ (or $G_{2}$) divides further unless $G_{2}$ (or $G_{1}$) is already dispersed. This keeps the total number of groups bounded by 2. As the settled agents keep the information required for the DFS traversal of all the groups, it helps the agents store the information using $O(\log n)$ memory. As we have different algorithms for the groups and different criteria for the groups to be divided, while dividing a group into two, each new group is given a unique label so that $G_{1}\,(G_{2})$ can execute their respective algorithms.

### v2

@@ 2 Impossibility Results and Lower Bound
(Impossibility without global communication) It is impossible to solve $k$-Dispersion on TVG for any $n\geq 6$ with the agents having $1$-hop visibility but no global communication even if agents are initially co-located. This impossibility result also holds if agents have infinite memory and knowledge of $n$.
@@
We prove this impossibility on TVG ring $\mathcal{R}$. Let $k=pn$ many agents be co-located at some node $v$ of ring $\mathcal{R}$, where $p$ is any positive integer. If the agents have 1-hop visibility, then as per our visibility model, the agent can sense the missing edge corresponding to some port $p$. There is no global communication. Therefore, the movement of the agents depends on 1-hop visibility and local communication. Since the algorithm is deterministic, the adversary can precompute the agents computation and accordingly remove an edge from $G$. Let the agents follow algorithm $\mathcal{A}$ to achieve dispersion. In a given round, if the movements of agents do not lead to dispersion, the adversary does not need to do anything. If the movements of agents lead to dispersion, we discuss how the adversary intervenes. Assume that at the end of round $r$, the agents successfully achieve dispersion. It means that the dispersion is not achieved at the beginning of round $r$, and after the CCM cycle at round $r$, the agents are in the dispersed configuration. Since $k=pn$, each node of $\mathcal{R}$ should have $p$ many agents in the dispersed configuration. At the beginning of round $r$, there is a node $u$ which has at most $p-1$ agents. Let $u_{1}$ and $u_{2}$ be two neighbours of node $u$ in ring $\mathcal{R}$. Let $z$ ($\leq p-1$) agents be at node $u$. If the adversary does not delete any edge at the beginning of round $r$, then the following is the computation of the agents according to $\mathcal{A}$: let $x_{1}$ agents move from node $u$ to $u_{1}$, $y_{1}$ agents move from node $u$ to $u_{2}$, $x_{2}$ agents move from node $u_{1}$ to $u$, and $y_{2}$ agents move from node $u_{2}$ to $u$. And accordingly, at the end of round $r$, there are $z+(x_{2}-x_{1})+(y_{2}-y_{1})=p$ agents at node $u$. The adversary knows the precomputed values of $x_{1},x_{2},y_{1},y_{2}$ and accordingly deletes an edge. A high-level idea behind the impossibility proof is given in Figure 1. We divide the cases as follows.
  • Case $\bm{x_{2}=0}$ or $\bm{y_{2}=0}$: W.l.o.g. let $x_{2}=0$. In this case, the adversary deletes the edge $(u,\,u_{2})$ at the beginning of round $r$. Therefore, the 1-hop view of agents at node $u_{1}$ does not change. Also, since global communication is not there, agents at node $u_{1}$ can not get the information about the missing edge $(u,\,u_{2})$ and hence the movement of agents at node $u_{1}$ will remain the same in round $r$. But the 1-hop view of agents at node $u$ and $u_{2}$ changes at round $r$. Therefore, at the end of round $r$, there are $z-x_{1}^{\prime}<p$ agents at the end of round $r$, where $x_{1}^{\prime}$ is the number of agents moving from node $u$ to $u_{1}$ due to the computation at $u$ based on the removed edge. Therefore, at the end of round $r$, the dispersion configuration is not achieved.
  • Case $\bm{x_{1}=0}$ or $\bm{y_{1}=0}$: W.l.o.g. let $x_{1}=0$. At the end of round $r$, there are $z+x_{2}+(y_{2}-y_{1})=p$ agents at node $u$. It is important to note that $x_{2}>0$ and $y_{2}>0$. Based on the values of $y_{1},\,y_{2}$, we have the following cases.
    $\rightarrow$ Sub-case $\bm{y_{1}=y_{2}:}$ The adversary deletes the edge $(u,\,u_{1})$. Therefore, the 1-hop view of agents at node $u_{2}$ does not change, but the 1-hop view of agents at node $u$ and $u_{1}$ changes at round $r$. Since global communication is not present, the movement of agents at node $u_{2}$ will remain the same in round $r$. In this case, from node $u$, more than $y_{1}$ agents or less than $y_{1}$ agents can not move to node $u_{2}$; otherwise, node $u_{2}$ contains either less than $p$ agents or greater than $p$ agents. If exactly $y_{1}$ agents move to node $u_{2}$, then the number of agents at node $u$ is $z$, and $z$ is less than $p$. Therefore, the dispersion configuration is not achieved at the end of round $r$.
    $\rightarrow$ Sub-case $\bm{y_{1}\neq y_{2}:}$ The adversary deletes the edge $(u,\,u_{2})$. Therefore, the 1-hop view of agents at node $u_{1}$ does not change, but the 1-hop view of agents at node $u$ and $u_{2}$ changes at round $r$. Since global communication is not present, the movement of agents at node $u_{1}$ will remain the same in round $r$. In this case, if agents move from node $u$ to $u_{1}$, then there are at least $p+1$ agents at node $u_{1}$ due to the view of node $u_{1}$. Therefore, the number of agents at node $u$ is $z+x_{2}$. If $z+x_{2}=p$, then $z+x_{2}=z+x_{2}+(y_{2}-y_{1})\implies y_{1}=y_{2}$. This gives a contradiction. Therefore, node $u$ does not contain $p$ agents at the end of round $r$. Therefore, the dispersion is not achieved in this case as well at the end of round $r$.
  • Case $\bm{x_{1},\,x_{2},\,y_{1},\,y_{2}>0}$: It is important to note that $(x_{2}-x_{1})\leq 0$ and $(y_{2}-y_{1})\leq 0$ are not possible. Therefore, either $(x_{2}-x_{1})>0$ or $(y_{2}-y_{1})>0$. W.l.o.g., let $(y_{2}-y_{1})>0$. In this case, the adversary deletes the edge $(u,\,u_{2})$ at the beginning of round $r$. Therefore, the 1-hop view of agents at node $u_{1}$ does not change, but the 1-hop view of agents at node $u$ and $u_{2}$ changes at round $r$. Since global communication is not present, the movement of agents at node $u_{1}$ will remain the same in round $r$. In this case, if more than $x_{1}$ or less than $x_{1}$ agents move from $u$ to $u_{1}$, then at node $u_{1}$, the number of agents is not $p$ due to the fact the view of node $u_{1}$. Therefore, the number of agents at node $u$ is $z+(x_{2}-x_{1})$. If $z+(x_{2}-x_{1})=p$, then $z+(x_{2}-x_{1})=z+(x_{2}-x_{1})+(y_{2}-y_{1})\implies(y_{2}-y_{1})=0$. This gives a contradiction. Therefore, at the end of round $r$, the number of agents at node $u$ is not $p$. Hence, the dispersion is not achieved at the end of round $r$.
@@
Figure 2: (a) If the adversary does not delete any edge at the beginning of round $r$, then this shows the computation of the agents. (b) If the adversary deletes the edge $(u,u_{2})$ at the beginning of round $r$, then the views of the agents at nodes $u_{1}$, $u_{2}$, and $u_{3}$ remain unchanged because they cannot detect the missing edge. In this situation, the agents at nodes $u$, $u_{1}$ and $u_{2}$ will attempt to send the same number of agents, but their movement will not be successful in round $r$. This fact makes the proof of Theorem 2.2 easy. Based on the cases mentioned in Theorem 2.2, the adversary decides which edge to delete at the start of round $r$. This outlines the high-level idea behind the proof of Theorem 2.2.
@@
(Impossibility without 1-hop visibility) It is impossible to solve $k$-Dispersion on TVG for any $n\geq 6$ considering global communication but without $1$-hop visibility even if agents are initially co-located. This impossibility result also holds if agents have infinite memory and knowledge of $n$.
@@
We prove this impossibility on TVG ring $\mathcal{R}$. Let $k=pn$ many agents be co-located at some node $v$ of ring $\mathcal{R}$, where $p$ is any positive integer. If the agents have zero-hop visibility, then the choice of movement of each agent depends only on the information present at the current node and the information attained via global communication. Since the algorithm is deterministic, the adversary can remove an edge from $G$ based on the movement of the agents. Let the agents follow algorithm $\mathcal{A}$ to achieve dispersion. In a given round, if the movement of agents does not lead to dispersion, the adversary does not need to do anything. If the movement of agents leads to dispersion, the adversary has to intervene. Assume that at the end of round $r$, the agents successfully achieve dispersion. It means that the dispersion is not achieved at the beginning of round $r$, and after the CCM cycle at round $r$, the agents are in the dispersed configuration. Since $k=pn$, each node of $\mathcal{R}$ should have $p$ many agents in the dispersed configuration. At the beginning of round $r$, there is a node $u$ which has at least $p+1$ many agents, and there is a node $v$ which has at most $p-1$ agents. Let $u_{1}$ and $u_{2}$ be two neighbours of node $u$ in ring $\mathcal{R}$. Based on the agent’s movement from node $u$, we give the strategy for the adversary at the beginning of round $r$ such that the dispersed configuration is not achieved at the end of round $r$. A high-level idea behind the impossibility proof is given in Figure 2. We divide the cases as follows.
  • Case 1. According to $\mathcal{A}$, if all the agents at node $u$ stay after $CCM$, then the adversary does not delete any edge at the beginning of round $r$. Since the agents of node $u$ do not move, therefore, at node $u$, there are at least $p+1$ agents at the end of round $r$. Hence, the dispersed configuration is not achieved at the end of round $r$.
  • Case 2. According to $\mathcal{A}$, at least one agent at node $u$ move to $u_{1}$ or $u_{2}$ after $CCM$ at round $r$. W.l.o.g., at least one agent at node $u$ moves to node $u_{1}$, and no agents move from $u$ to $u_{2}$. Let $x$ agents move to node $u_{1}$. The adversary deletes edge $(u,\,u_{1})$ at the beginning of round $r$. Since agents have 0-hop visibility, they can not understand the deletion of the edge at round $r$. Since edge $(u,u_{1})$ has been deleted, the movement of agents will be unsuccessful. In this case, at the end of round $r$, at node $u$, there will be at least $p+1$ agents. Hence, the dispersed configuration is not achieved at the end of round $r$.
  • Case 3. Let $z$ agents be at node $u$ at the beginning of round $r$, where $z>p$. According to $\mathcal{A}$ at round $r$, at least one agent at node $u$ moves to $u_{1}$, and at least one agent at node $u$ moves to node $u_{2}$. Let $x_{1}$ agents from node $u$ move to node $u_{1}$, and $y_{1}$ agents from node $u$ move to node $u_{2}$. Note that $x_{1},\,y_{1}\geq 1$. Let $x_{2}$ agents from node $u_{1}$ move to node $u$, and $y_{2}$ agents from node $u_{2}$ move to node $u$, where $x_{2},\,y_{2}\geq 0$. If the adversary does not delete any edge, then at the end of round $r$, there are $(z-x_{1}-y_{1}+x_{2}+y_{2})=p$ agents at node $u$. The adversary knows the precomputed values of $x_{1},x_{2},y_{1},y_{2}$ and accordingly deletes an edge. In this case, we divide this case into two sub-cases.
    • Sub-case $\bm{x_{1}=x_{2}}$ and $\bm{y_{1}=y_{2}:}$ At the end of round $r$, there are $z$ agents at node $u$, and $z>p$. Therefore, the dispersion configuration is not achieved at the end of round $r$.
    • Sub-case $\bm{x_{1}\neq x_{2}}$ or $\bm{y_{1}\neq y_{2}:}$ W.l.o.g., let $x_{1}\neq x_{2}$. In this case, the adversary deletes the edge $(u,\,u_{1})$ at the beginning of round $r$. Therefore, the movement of agents from node $u$ to $u_{1}$ and node $u_{1}$ to $u$ becomes unsuccessful. Since agents are unaware of the deletion of edge, the movement of agents from node $u$ to $u_{2}$ and node $u_{2}$ to $u$ becomes successful. Therefore, the number of agents at node $u$ is $z-y_{1}+y_{2}$. To achieve the dispersed configuration at the end of round $r$, the value of $z-y_{1}+y_{2}$ must be $p$. Therefore, $z-y_{1}+y_{2}=z-x_{1}-y_{1}+x_{2}+y_{2}\implies x_{1}=x_{2}$, and it is a contradiction due to $x_{1}\neq x_{2}$. Therefore, the value of $z-y_{1}+y_{2}$ is either less than $p$ or greater than $p$. Hence, the dispersion configuration is not achieved at the end of round $r$.
This completes the proof. ∎
Remark 1
It is important to note that the proof of both impossibility results is valid when $k=pn$, where $p$ is some constant, i.e., agents need 1-hop visibility and global communication to solve $pn$-Dispersion on TVG. In Section 3, we provide an algorithm to solve $(n+1)$-Dispersion on TVG with local communication and zero-hop visibility, and extend this idea in Section 4.1 to solve $(pn+1)$-Dispersion on TVG. The aforementioned proofs do not hold when $k\in[pn+2,\,(p+1)(n-1)]$, where $p\in\mathbb{N}$. Therefore, it is an open question whether 1-hop visibility and global communication are necessary to solve $k$-Dispersion on TVG for $k\in[pn+2,\,(p+1)(n-1)]$, where $p\in\mathbb{N}$. Also, it remains open to come up with the necessary conditions to solve $k$-Dispersion on TVG for $k<n$ given the algorithm of [16] that works with global communication and 1-hop visibility is a valid solution.
Remark 2
Note that both proofs are valid in the dynamic graph model mentioned in [16] since TVG is a stronger graph model. Therefore, agents need 1-hop visibility and global communication to solve the dispersion of $pn$ agents in the dynamic graph model mentioned in [16] as well.
Now, we provide the memory lower bound result to solve $k$-Dispersion on TVG. The proof of this lemma is motivated from [3]. In [3], they have provided the memory lower bound $\Omega(\log k)$ to solve dispersion on a static graph when $k\leq n$. We extend their idea for the memory lower bound when $k>n$.
Lemma 1
Assuming all agents are given the same amount of memory, agents require $\Omega(\log n)$ bits of memory each for any deterministic algorithm to solve $k$-Dispersion on TVG when $k\geq n+1$, and this proof is valid even if the adversary does not delete any edge at any round from footprint $G$.
Proof
Suppose all agents have $o(\log n)$ bits of memory. Each agent’s state space is $2^{o(\log n)}=n^{o(1)}$. Since there are at least $n+1$ agents, by pigeonhole principle, there exist at least $k-n+2$ agents with the same state space. Let agents $a_{1}$, $a_{2}$, …, $a_{k-n+2}$ have the same state space. Let us suppose that all agents are initially co-located at some node. Since all agents run the same deterministic algorithm, and agents $a_{1}$, $a_{2}$, …, $a_{k-n+2}$ are co-located initially, they will run the same moves. It means that agents $a_{1}$, $a_{2}$, …, $a_{k-n+2}$ remain stay together. Therefore, by the end of the algorithm, agents $a_{1}$, $a_{2}$, …, $a_{k-n+2}$ are at the same node. Solving dispersion requires a configuration where there is at least one agent at every node. We will never solve dispersion because there is no way for agents $a_{1}$, $a_{2}$, …, $a_{k-n+2}$ to settle down on at least two different nodes. Therefore, all $k$ agents are at most $n-1$ nodes by the end of the algorithm. Thus, to solve the dispersion problem, agents need $\Omega(\log n)$ memory each. This proof is for static graphs, and $TVG$ is a weaker model than static. Therefore, the same proof is also valid for TVGs. ∎
@@ 3 The Dispersion on General Graphs
In this section, we introduce an algorithm that solves the $(n+1)$-Dispersion on TVG when $n+1$ agents are co-located at some node in a general graph, and agents do not have global communication and 1-hop visibility, i.e., we assume that agents are equipped with zero-hop visibility and local communication. We use the idea of depth-first search (DFS) traversal by mobile agents that also solves the dispersion on static graphs [3]. For the sake of completeness, let us recall the idea of DFS traversal by a group of agents. At each node $v$, a group of agents chooses an unexplored port $p$ and traverses through it: (i) if it reaches a node $u$ which is an empty node (agents can understand this by the fact that there is no agent settled at $u$), then one agent settles at $u$, and stores port $p^{\prime}$ (through which it entered node $u$) as its parent port. Now, the unsettled agents try to move through port $(p^{\prime}+1)$ mod $\delta_{u}$ in the explore state, (ii) else, if the group of agents reach a node which already has a settled agent, the agent returns to node $v$ in the backtrack state. If the unsettled agents reach node $v$ in the backtrack state, then the agents compute $q=(p+1)$ mod $\delta_{v}$. If port $q$ is unexplored, it moves through port $q$ in the explore state. Else, if port $q$ equals the parent port of $v$, then the agents move through port $q$ in the backtrack state. This algorithm takes $4m$ rounds to run on any graph, where $m$ is the number of edges in that graph. The information required for the group of agents at each node to run this algorithm can be stored by the settled agent that is already present at that node.
@@
High-level idea: No group of agents moves indefinitely through which its movement was unsuccessful. Initially, there is only one group as the configuration is rooted. Whenever the group has an unsuccessful move, the agents divide into two groups, namely $G_{1}$ and $G_{2}$. These two groups run their dispersion algorithm via DFS separately. A missing edge can make the movement unsuccessful for both groups. On unsuccessful movement, $G_{1}$ never deviates from its DFS traversal path, whereas $G_{2}$ deviates after some finite waiting period. This keeps the algorithm running, and either $G_{2}$ gets dispersed, or $G_{1}$ can move one hop according to its DFS traversal. If $G_{2}$ gets dispersed, $G_{1}$ can understand it within some finite rounds and gets divided into two groups further. Else, if $G_{1}$ gets dispersed by moving one hop again and again, $G_{2}$ can understand it within some finite rounds and gets divided into two groups further. We ensure that none of $G_{1}$ (or $G_{2}$) divides further unless $G_{2}$ (or $G_{1}$) is already dispersed. This keeps the total number of groups bounded by 2. As the settled agents keep the information required for the DFS traversal of all the groups, it helps the agents store the information using $O(\log n)$ memory. As we have different algorithms for the groups and different criteria for the groups to be divided, while dividing a group into two, each new group is given a unique label so that $G_{1}\,(G_{2})$ can execute their respective algorithms.